# PEFTGuard: Detecting Backdoor Attacks Against Parameter-Efficient Fine-Tuning

### v1

@@ Introduction
Transfer Across Different Attacks. We aim to explore whether $\mathsf{PEFTGuard}$ possesses Zero-Shot Transferability from known attacks (those the model has been trained on) to the detection of unknown attacks (those the model has not encountered), using adapters trained on four different attack methods on AG News as shown in Table 16. Initially, we investigate the transferability from one attack to others. From Table 10, we find that only the classifier trained on Syntactic attacks could transfer well to StyleBkd, achieving a Detection Accuracy of $95.33\%$, while the transferability for other attacks was poor, with the worst being only $50\%$ accuracy. However, we also discover that despite the poor accuracy, the AUC indicated that $\mathsf{PEFTGuard}$ possesses some potential for zero-shot transfer.
To enhance the zero-shot transferability across different attacks, we employ Supervised Contrastive Loss for contrastive learning [30], training on three different attack datasets. Then, by combining classifiers through model fusion (using parameter averaging across three models trained on the same dataset), we successfully develop the detection capabilities for unknown attacks. As shown in Table 11, through contrastive learning and model fusion, $\mathsf{PEFTGuard}$, after being trained on any three attack datasets, demonstrates better zero-shot detection capabilities on unknown attacks. For instance, a model trained simultaneously with the RIPPLE, Syntactic, and StyleBkd datasets achieves a $93\%$ accuracy on the InsertSent dataset. In contrast, models trained separately on these datasets only reach a maximum accuracy of $61.33\%$ when transferred to InsertSent.
@@
Adaptive Attacking Scenarios. Therefore, we discuss the adversary’s use of different perturbations to attack the $\mathsf{PEFTGuard}$ classifier, including the addition of Gaussian noise, FGSM, I-FGSM, PGD, and C&W methods, with the adapters trained on IMDB based on the Llama-2-7B model with LoRA. As shown in Section 6.4, we consider two methods of adding Gaussian noise. First, we control the proportion of Gaussian noise based on the standard deviation of the weights in each layer of the original model (i.e., scaled by standard deviation). Secondly, we set the standard deviation of the noise equal to the standard deviation of the model’s original weights (scale = $5$), and adjust the proportion of Gaussian noise added within the total parameters of the model (i.e., Proportional to Parameter Size). For optimization-based adversarial attacks, we assume that the adversary can optimize the adversarial examples based on their trained model and transfer them to our $\mathsf{PEFTGuard}$ model. Note that here we consider a strong adversary that can train another detection model using the same training dataset and hyperparameters as $\mathsf{PEFTGuard}$.
Results Analysis. As shown in Section 6.4, as the proportion of noise increases, both methods affect the performance of the backdoored model, reducing both CA and ASR, but the ASR on the $\mathsf{PEFTGuard}$ classifier remains at $0\%$, indicating that $\mathsf{PEFTGuard}$ is robust against Gaussian noise. For FGSM, when $\epsilon=1\times 10^{-3}$, although the ASR on $\mathsf{PEFTGuard}$ can reach $92\%$, the CA of the backdoored model drops from $97.1\%$ to $44.9\%$, and the ASR also decreases from 0.999 to 0.139. The same conclusion can be seen in I-FGSM, PGD, and C&W. We also observe that when the intensity of the attack increases, the ASR on $\mathsf{PEFTGuard}$ can rise to $100\%$ in the first two and $86.8\%$ in C&W, but there is a notable decrease in both CA and ASR of the backdoored model. This indicates that even if the adversary can adjust the adapter’s weights to evade detection by $\mathsf{PEFTGuard}$ through training their own classifier, such adjustments significantly weaken the overall performance of the model, representing a considerable cost to the adversary.
@@
Although $\mathsf{PEFTGuard}$ performs well in various scenarios and exhibits some generalizability, there are several limitations.
  • For LLM models with different architectures, we design different network structures in $\mathsf{PEFTGuard}$ because transferring across different model structures involves changes in several dimensions (e.g., number of self-attention modules and dimension of projection matrices), which require redesigning the network structure to adapt to.
  • Although we have analyzed some aspects of model input, we do not clearly explore whether there is a characteristic change between PEFT fine-tuning and full-model fine-tuning backdoors. Also, it is unclear whether $\mathsf{PEFTGuard}$ is applicable for detecting backdoors in fully trained models (considering that our input is the product of multiplied adapters, i.e., the incremental part of the model’s trained parameters). We leave it as our future work.
  • $\mathsf{PEFTGuard}$ currently requires white-box access to the parameters of the adapters, making it unsuitable for models that only provide API access.
@@
LLM. Recently, LLMs have achieved great success in the field of NLP [76, 45]. Trained on large amounts of text, LLMs have developed strong language modeling capabilities and assisted humans in solving complex tasks, such as OpenAI’s ChatGPT [47] and GPT-4 [46] and Microsoft’s Copilot systems [44]. There are many open-source LLMs (such as Llama-3 [43], Mixtral [29] and Qwen [3]), and it’s necessary to fine-tune models for specific downstream tasks.
@@
Backdoor Attacks. In the field of NLP, researchers have studied backdoor attacks. Inserting triggers into the data is the most common and effective attack method [33, 9, 69] leads models to malicious behaviors. More seriously, injecting backdoors into the training of LLMs can cause models to generate toxic responses when triggered, leading to severe consequences [62, 64, 66], and can also cause significant security issues in applications based on LLMs [68]. This raises concerns about open-source LLMs on the internet. In addition, due to the effectiveness of PEFT, many people share their PEFT adapter models online to accomplish various downstream tasks, which may result in the spread of backdoors [36]. Therefore, backdoor detection methods for PEFT are highly demanded.
Backdoor Detection on LLMs. Early methods for backdoor detection in NLP tasks include Trojan-Miner [2] targets DNN-based text classification tasks, using a seq-2-seq model to probe suspicious classifiers and generate sentences that may contain trojan triggers to detect the backdoors. AttenTD [40] is an attention-based detector that provides a set of neutral trigger candidates and distinguishes backdoored models through attention anomalies. PICCOLO [38] determines the presence of backdoors by analyzing the model’s sensitivity to trigger words. Note that one of the most similar methods as PEFTGuard is MNTD [65], which relies on representation vectors of model outputs, whereas $\mathsf{PEFTGuard}$ directly leverages the model parameters. However, with the rise of decoder-only structures in LLMs [4, 47], there is an increasing focus on tasks involving the generation of coherent content. The main behavior of backdoor attacks on this task is typically to induce the LLMs to generate wrong or toxic output, which brings new challenges to traditional detection methods. Specifically, trigger generation requires optimizing text triggers through changes in classification labels, and optimization-based trigger inversion methods face the problem of generating precise triggers from changes in the discrete output domain in generation tasks. Moreover, attention analysis methods that rely on preset trigger words face challenges because adding a new word can shift the model’s attention to fit the context and content, making detection much harder.
@@ Conclusion
In this paper, we conduct the first in-depth and comprehensive analysis to reveal the security vulnerabilities of backdoored PEFT-based adapters. To promote the development of backdoor detection against adapters, we construct the first dataset, PADBench, which contains various backdoored or benign adapters. Meanwhile, we propose the first backdoor detection method, named $\mathsf{PEFTGuard}$. $\mathsf{PEFTGuard}$ does not require any additional information during the detection. Instead, it only requires access to the parameters of the PEFT adapter. $\mathsf{PEFTGuard}$ achieves state-of-the-art performance in detecting various types of adapters, which are generated from different datasets, backdoor attack algorithms, PEFT methods, and various base models (including different model parameter sizes, both base and fine-tuned models, diverse model structures, and different attention mechanisms). In addition, $\mathsf{PEFTGuard}$ demonstrates a zero-shot transferability across different PEFT fine-tuning methods, adapter ranks, and backdoor attacks. Furthermore, we show that $\mathsf{PEFTGuard}$ is robust against different adaptive attacks. Overall, we hope that the proposed PADBench and $\mathsf{PEFTGuard}$ will play a key role in advancing the security governance of adapters in the open-source platforms.
Appendix A Ablation Study about Zero-shot Transferability on Unknown Attacks.
@@
In Section 6.4, we employ Contrastive Learning (CL), combined with model fusion via parameter averaging, enabling models trained on three different attack datasets to generalize to another unknown attack dataset. To investigate the effectiveness of each component, we conduct ablation studies, specifically targeting the unknown attack as Syntactic. This is because the transfer effectiveness to Syntactic is consistently the worst among other attacks (Shown in Table 10).
@@
Appendix B Related Work of Backdoor Mitigation Method.
The goal of backdoor mitigation methods is to directly eliminate backdoors from models, and these methods can be used in conjunction with detection techniques to first identify and then remove the backdoors. In this paper, we focus on methods to eliminate backdoors through training techniques [70, 54, 51, 75].
Experiments by Yao et al. [70] and Sha et al. [54] both indicate that fine-tuning backdoored models on a clean subset of training samples can mitigate the backdoors. Rafailov et al. [51] propose Direct Preference Optimization (DPO), an optimization method specifically designed for LLMs. This method utilizes the mapping relationship between reward functions and optimal policies, demonstrating that this constrained reward maximization problem can be accurately optimized through single-stage policy training. By setting texts that contain backdoor triggers but have normal answers as preferred texts, the DPO method can effectively eliminate backdoor influences in the model. Zhang et al. [75] propose the Fine-mixing method, which considers the clean pre-trained model weights before fine-tuning on clean data and mixes the backdoored weights with clean pre-trained weights. In addition, they utilize Embedding Purification (E-PUR) to detect and eliminate potential backdoor techniques within embeddings.
Table 15: Detailed parameter settings for the C&W attack.
Parameter Set Settings
$P_{3}$ $c=0.1$, $\kappa=0$, $\text{iter}=30$, $\text{lr}=1\times 10^{-4}$
$P_{4}$ $c=0.1$, $\kappa=5$, $\text{iter}=30$, $\text{lr}=1\times 10^{-4}$
$P_{5}$ $c=0.5$, $\kappa=10$, $\text{iter}=30$, $\text{lr}=1\times 10^{-4}$
Figure 7: Architecture of $\mathsf{PEFTGuard}_{\rm cls}$.
Figure 8: Alternative Architectures of $\mathsf{PEFTGuard}$, representing two examples among many potential configurations.
Table 16: Details of our PADBench Datasets. “-” indicates empty because the instruction-following datasets can’t be evaluated for clean accuracy. “*” indicates that in IMDB and AG News datasets, the benign adapters remain the same across different attacks in the same training dataset.(Unless specified otherwise, the target projection matrices in the PEFT method are, by default, applied to the query and value matrices.)
@@
ViT-base CIFAR-10 BadNet LoRA $16$ $/250$ 250 $0.985$ $0.984$ $0.921$
Table 17: Overview of Backdoor Injection Datasets and Tasks

### v2

@@ Introduction
Transfer Across Different Attacks. We aim to explore whether $\mathsf{PEFTGuard}$ exhibits Zero-Shot Transferability from known attacks (those the model has been trained on) to the detection of unknown attacks (those the model has not encountered), using adapters trained on four different attack methods on AG News as shown in Table 15. Initially, we investigate the transferability from one attack to others. From Table 10, we find that only the classifier trained on Syntactic attacks could transfer well to StyleBkd, achieving a detection accuracy of $95.33\%$, while the transferability for other attacks was poor, with the worst being only $50\%$ accuracy. However, we also discover that despite the poor accuracy, the AUC indicated that $\mathsf{PEFTGuard}$ possesses some potential for zero-shot transfer.
To enhance the zero-shot transferability across different attacks, we employ supervised contrastive loss for contrastive learning [32], training on three different attack datasets. Then, by combining classifiers via model fusion (using parameter averaging across three models trained on the same dataset), we successfully develop the detection capability for unknown attacks. As shown in Table 11, by leveraging contrastive learning and model fusion, $\mathsf{PEFTGuard}$, after being trained on any three attack datasets, demonstrates better zero-shot detection capability on unknown attacks. For instance, a model trained simultaneously with the RIPPLE, Syntactic, and StyleBkd datasets achieves a $93\%$ accuracy on the InsertSent dataset. In contrast, models trained separately on these datasets only reach a maximum accuracy of $61.33\%$ when transferred to InsertSent.
@@
Adaptive Attacking Scenarios. Therefore, we discuss the adversary’s use of different perturbations to attack $\mathsf{PEFTGuard}$ classifier, including the addition of Gaussian noise, FGSM, I-FGSM, PGD, and C&W methods, with the adapters trained on IMDB based on the Llama-2-7B model with LoRA. As shown in Section 6.4, we consider two methods of adding Gaussian noise. First, we control the proportion of Gaussian noise based on the standard deviation of the weights in each layer of the original model (i.e., scaled by standard deviation). Secondly, we set the standard deviation of the noise equal to the standard deviation of the model’s original weights (scale = $5$), and adjust the proportion of Gaussian noise added within the total parameters of the model (i.e., proportional to parameter size). For optimization-based adversarial attacks, we assume that the adversary can optimize the adversarial examples based on their trained model and transfer them to our $\mathsf{PEFTGuard}$ model. Note that here we consider a strong adversary that can train another detection model using the same training dataset and hyperparameters as $\mathsf{PEFTGuard}$.
Results Analysis. As shown in Section 6.4, as the proportion of noise increases, both methods affect the performance of the backdoored model, reducing both CA and ASR, but the ASR on $\mathsf{PEFTGuard}$ classifier remains at $0\%$, indicating that $\mathsf{PEFTGuard}$ is robust against Gaussian noise. For FGSM, when $\epsilon=1\times 10^{-3}$, although the ASR on $\mathsf{PEFTGuard}$ can reach $92\%$, the CA of the backdoored model drops from $97.1\%$ to $44.9\%$, and the ASR also decreases from 0.999 to 0.139. The same conclusion can be seen in I-FGSM, PGD, and C&W. We also observe that when the intensity of the attack increases, the ASR on $\mathsf{PEFTGuard}$ can rise to $100\%$ in I-FGSM and PGD and $86.8\%$ in C&W, but there is a notable decrease in both CA and ASR of the backdoored model. This suggests that even if the adversary can adjust the adapter’s weights to evade detection by $\mathsf{PEFTGuard}$ by training their own classifier, such adjustments significantly degrade the overall performance of the model, representing a considerable cost to the adversary.
@@
Transferability. Since $\mathsf{PEFTGuard}$ relies on training a meta-classifier to detect backdoored adapters, any variation in input dimensions requires training a new classifier. This constraint limits the zero-shot transferability of $\mathsf{PEFTGuard}$ across different LLMs. Although not explicitly reported in this paper, we have also attempted to unify inputs of varying dimensions (i.e., different LLMs) by downsampling. Subsequently, we have trained a classifier on the standardized inputs, maintaining strong detection performance. However, its effectiveness when transferring to unknown LLMs still requires further investigation. Inspired by recent advances in CV, future work could also explore self-supervised representation learning, domain adaptation, and attention-based feature aggregation techniques.
Practical Deployment. Although $\mathsf{PEFTGuard}$ is effective, when applied to a completely new detection scenario, it requires substantial time for training and deployment. However, as shown in Figure 6, we can achieve strong detection performance with only $100$ training samples, which can reduce deployment time. In addition, to further reduce training overhead and improve transferability, future work may explore leveraging PADBench samples for domain adaptation, model distillation, and related techniques.
Backdoor Pattern Explanation. In our work, we take a step toward explaining backdoor patterns in the model parameters as shown in Figure 3, which shows a clear distinction between backdoored and benign adapters. Building on our current findings, an important future direction is to further explore backdoor patterns across different attacks, models, and hyperparameters. In particular, analyzing these patterns in the task vector space could be promising, and prior works in the CV domain [47, 36] may provide helpful insights.
@@
LLM. Recently, LLMs have achieved great success in NLP domain [86, 50]. Trained on large amounts of text, LLMs have developed strong language modeling capabilities and assisted humans in solving complex tasks, such as OpenAI’s ChatGPT [52] and GPT-4 [51] and Microsoft’s Copilot systems [49]. There are many open-source LLMs (such as Llama-3 [48], Mixtral [31] and Qwen [3]), and it’s necessary to fine-tune models for specific downstream tasks.
@@
Backdoor Attacks. In the field of NLP, researchers have studied various backdoor attacks. Inserting triggers into the data is the most common and effective attack method [35, 11, 74] leads models to malicious behaviors. More seriously, injecting backdoors into the training of LLMs can cause models to generate toxic responses when triggered, leading to severe consequences [66, 83, 71], and can also cause significant security issues in applications based on LLMs [73]. This raises concerns about open-source LLMs on the internet. In addition, due to the effectiveness of PEFT, many people share their PEFT adapter models online to accomplish various downstream tasks, which may result in the spread of backdoors [39]. Besides, Dong et al. [14] also propose two novel backdoor attacks targeting adapters, POLISHED and FUSION, which successfully manipulate the LLMs to perform malicious actions. Therefore, backdoor detection methods for PEFT are highly demanded.
Backdoor Detection on LLMs. Early methods for backdoor detection in NLP tasks include Trojan-Miner [2] targets DNN-based text classification tasks, using a seq-to-seq model to probe suspicious classifiers and generate sentences that may contain trojan triggers to detect the backdoors. AttenTD [44] is an attention-based detector that provides a set of neutral trigger candidates and distinguishes backdoored models through attention anomalies. PICCOLO [41] determines the presence of backdoors by analyzing the model’s sensitivity to trigger words. Note that one of the most similar methods as PEFTGuard is MNTD [70], which relies on representation vectors of model outputs, whereas $\mathsf{PEFTGuard}$ directly leverages the model parameters. Besides, Zeng et al. [78] propose the first framework, CLIBE, for detecting dynamic backdoors in transformer-based NLP models by introducing few-shot perturbations into the suspect model’s parameters. However, with the rise of decoder-only architectures in LLMs [4, 52], there is an increasing focus on tasks involving the generation of coherent content. The main behavior of backdoor attacks on this task is typically to induce the LLMs to generate incorrect or toxic text outputs, which brings new challenges to traditional detection methods. Specifically, trigger generation requires optimizing text triggers through changes in classification labels, and optimization-based trigger inversion methods struggle to generate precise triggers from changes in the discrete output domain in generation tasks. Moreover, attention analysis methods that rely on preset trigger words face challenges because adding a new word can shift the model’s attention to fit the context, making detection much more challenging.
@@ Conclusion
In this paper, we conduct the first in-depth and comprehensive analysis to reveal the security vulnerabilities of backdoored PEFT-based adapters. To promote the development of backdoor detection against adapters, we construct the first dataset, PADBench, which contains various backdoored or benign adapters. Meanwhile, we propose the first backdoor detection method, named $\mathsf{PEFTGuard}$. It does not require any additional information during the detection, only access to the PEFT adapter parameters. $\mathsf{PEFTGuard}$ achieves state-of-the-art performance in detecting various types of adapters, which are generated from different datasets, backdoor attacks, PEFT methods, and various base LLMs. In addition, $\mathsf{PEFTGuard}$ demonstrates a zero-shot transferability across different PEFT methods, adapter ranks, and backdoor attacks. Furthermore, we show that $\mathsf{PEFTGuard}$ is robust against different adaptive attacks. Overall, we hope that the proposed PADBench and $\mathsf{PEFTGuard}$ will play a key role in advancing the security governance of adapters in the open-source platforms.
Acknowledgments
We sincerely thank the reviewers and our shepherd for their constructive feedback, which improved the quality of this work. This work is supported by the National Natural Science Foundation of China (No.62425205, No.62376210, and No.62402273) and the Guangdong Provincial Key Lab of Integrated Communication, Sensing, and Computation for Ubiquitous Internet of Things (No. 2023B1212010007).
Appendix A Ablation Study about Zero-shot Transferability on Unknown Attacks
@@
In Section 6.4, we employ Contrastive Learning (CL), combined with model fusion via parameter averaging, enabling models trained on three different attack datasets to generalize to another unknown attack dataset. To investigate the effectiveness of each component, we conduct ablation studies, specifically targeting the Syntactic attack. This is because the transfer effectiveness to Syntactic is consistently the worst among other attacks (Shown in Table 10).
@@
Appendix B Related Work on Backdoor Mitigation Methods
Backdoor mitigation methods aim to directly eliminate backdoors from models and can be combined with detection techniques to first identify and then remove them. We focus on methods to eliminate backdoors through training techniques [76, 59, 56, 85]. Experiments by Yao et al. [76] and Sha et al. [59] both indicate that fine-tuning backdoored models on a clean subset of training samples can mitigate the backdoors. Rafailov et al. [56] propose Direct Preference Optimization (DPO), an optimization method specifically designed for LLMs. This method utilizes the mapping relationship between reward functions and optimal policies, demonstrating that this constrained reward maximization problem can be accurately optimized through single-stage policy training. By setting texts that contain backdoor triggers but have normal answers as preferred texts, the DPO method can effectively eliminate backdoor influences in the model. Zhang et al. [85] propose the Fine-mixing method, which considers the clean pre-trained model weights before fine-tuning on clean data and mixes the backdoored weights with clean pre-trained weights. In addition, they utilize Embedding Purification (E-PUR) to detect and eliminate potential backdoor techniques within embeddings.
(a) Architecture of $\mathsf{PEFTGuard}_{\rm cls}$.
(b) Alternative Architectures of $\mathsf{PEFTGuard}$, representing two examples among many potential configurations.
Figure 7: Illustration of $\mathsf{PEFTGuard}$ architectures.
Table 15: Details of our PADBench. “-” indicates empty because the instruction-following datasets can’t be evaluated for clean accuracy. “*” indicates that in IMDB and AG News datasets, the benign adapters remain the same across different attacks in the same training dataset.(Unless specified otherwise, the target projection matrices in the PEFT method are, by default, applied to the query and value matrices.)
@@ Appendix B Related Work on Backdoor Mitigation Methods
ViT-base CIFAR-10 BadNets LoRA $16$ $/250$ 250 $0.985$ $0.984$ $0.921$
Table 16: Summary of backdoor injection datasets and tasks.
@@
Table 17: Parameter settings for the C&W attack.
Parameter Set Settings
$P_{3}$ $c=0.1$, $\kappa=0$, $\text{iter}=30$, $\text{lr}=1\times 10^{-4}$
$P_{4}$ $c=0.1$, $\kappa=5$, $\text{iter}=30$, $\text{lr}=1\times 10^{-4}$
$P_{5}$ $c=0.5$, $\kappa=10$, $\text{iter}=30$, $\text{lr}=1\times 10^{-4}$